OpenStudy (anonymous):

How many solutions does a system of two linear equations have if the slope of each equation is the same, but the y-intercepts are different? infinitely many none one none of the above

OpenStudy (anonymous):

If the slopes are the same, then the lines are parallel, which means they never intersect and there are no solutions.

OpenStudy (anonymous):

do you think you can help me with my math I'm so bad at it and i need some help

OpenStudy (anonymous):

and so would i put none of the above

OpenStudy (anonymous):

depends on what you ask. I'll help where I can.

OpenStudy (anonymous):

Just none

OpenStudy (anonymous):

okay thank you and hold on

OpenStudy (anonymous):

Select the system of linear equations that has the solution (7, 0).

OpenStudy (anonymous):

A. x-y=7 B. x+y=9 & 7x+3y=10 C.-5x-y=-35 & x-y=7 D.x+y=7 &3x-4y=21 & 3x+4y=-20 this is the choices

OpenStudy (anonymous):

The solution only has 2 numbers, so there have to be 2 equations. Can't be A or D

OpenStudy (anonymous):

Plug in 7 for x and 0 for y to test the other systems.

OpenStudy (anonymous):

i got b for the answer

OpenStudy (anonymous):

The first equation in B is x + y = 9 7 + 0 is not 9, so that can't be the answer

OpenStudy (anonymous):

oh i see now

OpenStudy (anonymous):

and my last questions is how many solutions does x-3y=-8 & x+3y=8

OpenStudy (anonymous):

infinitely many none one none of the above

OpenStudy (anonymous):

The slope of the first equation is 1/3. The slope of the second equation is -1/3. Different slopes → 1 solution

OpenStudy (anonymous):

thank you so much are you good at graphs

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

I'm going to send you one hold on

OpenStudy (anonymous):

(0, -5) (1, 3) (0, 0) (-1, -3)

OpenStudy (anonymous):

The yellow box is kind of covering it, but the solution is the point where the blue lines cross.

OpenStudy (anonymous):

is it going to be -1,-3

OpenStudy (anonymous):

yeah that looks right

OpenStudy (anonymous):

(3, -6) (1, 0) (0, 4.5) (2, 7)

OpenStudy (anonymous):

same thing, pick where the lines cross

OpenStudy (anonymous):

is it 2,7

OpenStudy (anonymous):

no. 

OpenStudy (anonymous):

oh is it 3,-6

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

yaaay thank you so much with the help

OpenStudy (anonymous):

you're welcome